Jill has $3.50 in nickels and dimes. If she has 50 coins, how many of each type of coin does she have?
Answer by piglet162431(90)

soo
let statements:
let x be the number of nickels
let y be the number of dimes
since there are 50 coins, we can say
x+y=50
since in total the value of these coins are $3.50
.05x+.1y=3.50
to make things simply, why not times the equation by 100
5x+10y=350
Now use substitution method.
from the 1st equation, y=50-x
5x+10(50-x)=350
5x+500-10x=350
-5x=-150
x=30
y=20
There are 30 nickels and 20 dimes.